OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest whole number.

OpenStudy (anonymous):

possible answers 260 320 360 290

OpenStudy (anonymous):

y = kxn^2 this is a two step process: plug in values of y,x and n and solve for k then insert value of k into the above formula plug in x=5 and n=10 and solve for y

OpenStudy (anonymous):

260

OpenStudy (anonymous):

thanks ryan

OpenStudy (anonymous):

y=kxn^2 25 = k(3)(4)^2 k=25/48 then y=(25/48)(5)(10)^2 ~260.416...